OpenStudy (anonymous):

Please helpppp, fan, medal, testimony Segment AB has point A located at (4, 2). If the distance from A to B is 3 units, which of the following is the coordinate for point B? A (5, 2) B (1, 5) C (4, -1) D (-2, 2) I honestly think its C, but im not sure

OpenStudy (mathmath333):



OpenStudy (mathmath333):

u can form an equation of circle here \(\large \color{black}{\begin{align}(x-4)^2+(y-2)^2=3^2\hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (mathmath333):

from u can try to put each option in \((x,y)\) in LHS and see if it equals \(3^2\)

OpenStudy (anonymous):

I don't get it ): @mathmath333

OpenStudy (mathmath333):

\(\large \color{black}{\left. \begin{array}{l} A (5, 2) \\ B (1, 5) \\ C (4, -1)\\ D (-2, 2)\\ \end{array} \right\} =\text{4 options}}\) try to put each of them in equation of circle and see if it equals 3^2. if yes than that is the answer example.\((5,2)\) \(\large \color{black}{(x-4)^2+(y-2)^2=3^2\hspace{.33em}\\~\\ (5-4)^2+(2-2)^2\\ =1^2+0^2=1\neq 3\hspace{.33em}\\~\\ }\) so the first option is wrong like this try all other \(3\) options.

OpenStudy (anonymous):

so D (: @jdoe0001

OpenStudy (jdoe0001):

hmm

OpenStudy (jdoe0001):

well... in short, the point (4,2) is just a point and there's an infinite amount of lines that can go through it and any of those lines can be 3units long so the answer is an infinite amount HOWEVER you were given a few choices so, if you pick the choices as your 2nd point, and find the "slope" between those 2 points and it gives you 3, then that's one

OpenStudy (jdoe0001):

you could use the distance formula I gather, to check for the same thing the one choice that gives you a distance of 3

OpenStudy (jdoe0001):

well have you covered slopes? or the distance formula yet?

OpenStudy (mathmath333):

\(\large \color{black}{(x-4)^2+(y-2)^2=3^2\hspace{.33em}\\~\\ \normalsize \text{if u try option c here}\\~\\ (4-4)^2+(-1-2)^2\hspace{.33em}\\~\\ =0^2+(-3)^2=3^2\\~\\ so\\ LHS=RHS }\)

OpenStudy (anonymous):

@satellite73 hey you, youre my favorite on here. lol please help me understand

@ganeshie8 please help me):